OpenStudy (anonymous):

please help 14x+5/6=-2/6 solve for x

OpenStudy (mimi_x3):

Is it: \[\large \frac{14x+5}{6} =-\frac{2}{6} \]?

OpenStudy (anonymous):

... its only 5 over 6